OpenStudy (anonymous):

Find the next numbers in the pattern? 1/2, 1/4, 1/8, 1/16, … Choices: 1/32, 1/64, 1/148 1/24, 1/32, 1/40 1/32, 1/48, 1/64 1/32, 1/64, 1/128 I'm really confused? I just need help figuring out the Common Difference etc.

OpenStudy (anonymous):

the denominator is being multiplied by two each time

OpenStudy (anonymous):

You are doubling the last number, so 1/2, double the 2, so its 4, so 1/4, double 4 and get 8, and so on, get it?

OpenStudy (anonymous):

this is a geometric sequence so there is a common ratio it can be found by dividing the second term with the first term so the common ratio r=1/2 to find the next numbers, you have to multiply the previous number with the common ratio so the answer will be: 1/32, 1/64, 1/128

OpenStudy (yttrium):

\[a_{n} = \frac{ 1 }{ 2n }\] \[a_{1} = \frac{ 1 }{ 2(1) } = \frac{ 1 }{ 2 }\] \[a_{2} = \frac{ 1 }{ 2(2) } = \frac{ 1 }{ 4 }\] \[a_{3} = \frac{ 1 }{ 2(3) } = \frac{ 1 }{ 8 }\] Just continue and you will get the next of the sequence